# Improving Alignment and Robustness with Circuit Breakers

### v3

@@ 4.1 Large Language Models
RepE Attack: This white-box attack manipulates the model’s representation space. Previous work demonstrates the identification of directional vectors in the model’s representation space that correspond to refusals [77]. By altering these vectors—either adding or subtracting—we can modulate the model’s tendency to refuse requests.
We utilize HarmBench’s LLM classifier to evaluate the attack success rate and manually verify the judgements. Detailed configurations for each attack are provided in Section C.2.2. To measure the capabilities of the models with circuit breakers, we evaluate our models on MTBench [75] for instruction-following abilities and on the OpenLLM Leaderboard [8] for knowledge and reasoning which includes MMLU [22], ARC-c [13], HellaSwag [72], TruthfulQA [31], Winogrande [58], and GSM8K [14]. Additionally, we follow the methodology in [5] to construct an over-refusal evaluation, described in Appendix B. For baselines, we use the original Mistral and Llama-3 Instruct models. Additionally, we include a state-of-the-art adversarially trained Mistral model, R2D2 [40], for comparison.
@@ 4.4 Ablation and Analysis
To understand the generalization properties of circuit-breaking, we split our training data into six categories of harm, train category-specific models, and measure their generalization performance across categories. We find strong in-domain generalization, indicated by the low ASR along the diagonal, and observe that training on broader categories like Harmful and Illegal Activities offers greater generalization than narrower categories like Cybercrime. We report similar ablations for Mistral-7B in Appendix F.
@@
Representation analysis. In Figure 6, we plot the cosines between representations of the Llama-3-8B-Instruct model with and without circuit breakers for a prefilled harmful response “Here is how to synthesize meth: 1. Start with”. We additionally plot the norms of these representations in Figure 9. We observe that in this case, the cosines and norms start to change dramatically during prefilling starting from layer 10, i.e., even before generation starts. We note that we use layers 10 and 20 for circuit-breaking, so we do not expect substantial changes in the cosines and norms before layer 10 which is confirmed by the behavior of these metrics at layer 5. Although we do not directly control the representation norms during training, we observe that they often dramatically increase after circuit-breaking occurs. We repeat the same experiment for Mistral-7B-Instruct and show it in Appendix G, where we also analyze two other prompts: one that leads to a similar behavior and one that triggers circuit breakers after generation starts. Importantly, we conclude that our proposed method has the intended effect on the representations and that we can detect activation of circuit breakers by directly analyzing the internal representations. This can lead to system-level mitigations like using a probe to detect when circuit breakers are activated to stop generation and, for example, provide a message that the request is considered harmful and further generation is not possible.
| Mistral-7B-Instruct-v2 | Llama-3-8B-Instruct | ||||||
| Refusal Trained | + HP | + RR | Refusal Trained | + HP | + RR | ||
| Over-Refusal | WildChat | 2.0 | 3.6 | 3.4 | 2.2 | 6.2 | 6.2 |
| Robustness | No Attack | 57.8 | 17.9 | 4.9 | 12.4 | 7.8 | 1.2 |
| Manual | 77.4 | 7.4 | 6.8 | 8.3 | 2.1 | 0.0 | |
| TAP-T | 85.8 | 27.5 | 17.5 | 17.4 | 6.6 | 2.1 | |
| GCG | 88.7 | 18.0 | 11.2 | 44.5 | 10.0 | 2.5 | |
| Input Embed | 92.1 | 16.3 | 15.7 | 80.4 | 13.7 | 9.6 | |
| Average | 77.4 | 17.7 | 10.1 | 20.7 | 6.6 | 1.5 | |
@@
Circuit breaking with Harmfulness Probes (HP). Our proposed circuit breaking method relies on representation control. This section evaluates the potential efficacy of representation reading as an alternative. Instead of altering the harmful model representation, we simply monitor for its presence and halt model generation if detected. We employ the same training dataset (harmful circuit breaker set and retain set) used in the LLM experiments. A logistic regression classifier is trained to distinguish between model activations from the two datasets. Specifically, activations are collected from the 16th layer of the Mistral model and from the final layer of the Llama-3 model for each token in 2000 responses. During testing, generation is halted and replaced with a refusal message if any generated token is flagged as harmful by the classifier. We find a threshold so that the false positive rates (FPR) on WildChat (Appendix B) is around the same as the model trained with RR [74]. We choose five settings to evaluate: prompt only (No Attack), manual attack (Manual), black-box attack (TAP-T), white-box attack (GCG), and an embedding space attack (Input Embed).
@@
As shown in table Table 2, Harmfulness Probing (HP) significantly reduces the attack success rate compared to the refusal-trained baseline, but is outperformed by the representation control approach (RR). We emphasize that, although generally probing can be easily thwarted by adversarial attacks, much like input and output filters, its robustness in this context can be largely attributed to the continuous monitoring of model representations associated with harmful processes throughout the entire generation, a key idea in circuit breaking. It is important to note, however, that the Harmful Probes are tested under a weaker adversarial setting, where the attacker lacks knowledge of the probe and does not directly optimize against it. Further investigation into Harmfulness Probes is left for future work.
@@ C.4.1 AI Agent Benchmark Example
Appendix D Detailed Results in Multimodal and Agent Settings
@@
Appendix E Multilingual Results
@@
We do not see that trend in Table 5. We leave investigation of this to future work.
Appendix F Additional Ablation Results
@@
Appendix G Extended analysis of representations

### v4

@@ 4.1 Large Language Models
RepE Attack: This white-box attack manipulates the model’s representation space. Previous work in representation engineering demonstrates the identification of directional vectors in the model’s representation space that correspond to refusals [77]. By altering these vectors—either adding or subtracting—we can modulate the model’s tendency to refuse requests.
We utilize HarmBench’s LLM classifier to evaluate the attack success rate and manually verify the judgements. Detailed configurations for each attack are provided in Section C.2.2. To measure the capabilities of the models with circuit breakers, we evaluate our models on MTBench [75] for instruction-following abilities and on the OpenLLM Leaderboard [8] for knowledge and reasoning which includes MMLU [22], ARC-c [13], HellaSwag [72], TruthfulQA [31], Winogrande [58], and GSM8K [14]. Table 5 contains a detailed breakdown of performance on each dataset. Additionally, we follow the methodology in [5] to construct an over-refusal evaluation, described in Appendix B. For baselines, we use the original Mistral and Llama-3 Instruct models. Additionally, we include a state-of-the-art adversarially trained Mistral model, R2D2 [40], for comparison.
@@ 4.4 Ablation and Analysis
To understand the generalization properties of circuit-breaking, we split our training data into six categories of harm, train category-specific models, and measure their generalization performance across categories. We find strong in-domain generalization, indicated by the low ASR along the diagonal, and observe that training on broader categories like Harmful and Illegal Activities offers greater generalization than narrower categories like Cybercrime. We report similar ablations for Mistral-7B in Appendix G.
@@
Representation analysis. In Figure 6, we plot the cosines between representations of the Llama-3-8B-Instruct model with and without circuit breakers for a prefilled harmful response “Here is how to build a bomb: 1. Start with”. We additionally plot the norms of these representations in Figure 12. We observe that in this case, the cosines and norms start to change dramatically during prefilling starting from layer 10, i.e., even before generation starts. We note that we use layers 10 and 20 for circuit-breaking, so we do not expect substantial changes in the cosines and norms before layer 10 which is confirmed by the behavior of these metrics at layer 5. Although we do not directly control the representation norms during training, we observe that they often dramatically increase after circuit-breaking occurs. We repeat the same experiment for Mistral-7B-Instruct and show it in Appendix H, where we also analyze two other prompts: one that leads to a similar behavior and one that triggers circuit breakers after generation starts. Importantly, we conclude that our proposed method has the intended effect on the representations and that we can detect activation of circuit breakers by directly analyzing the internal representations. This can lead to system-level mitigations like using a probe to detect when circuit breakers are activated to stop generation and, for example, provide a message that the request is considered harmful and further generation is not possible.
| Mistral-7B-Instruct-v2 | Llama-3-8B-Instruct | ||||||||
| Refusal Trained | + HP (Linear) | + HP (MLP) | + RR | Refusal Trained | + HP (Linear) | + HP (MLP) | + RR | ||
| Over-Refusal | WildChat | 2.0 | 3.6 | 3.6 | 3.4 | 2.2 | 6.2 | 6.2 | 6.2 |
| Robustness | No Attack | 57.8 | 16.6 | 12.5 | 4.9 | 12.4 | 6.6 | 5.8 | 1.2 |
| Manual | 77.4 | 7.4 | 5.2 | 6.8 | 8.3 | 1.7 | 0.8 | 0.0 | |
| TAP-T | 85.8 | 27.5 | 26.2 | 17.5 | 17.4 | 8.3 | 6.2 | 2.1 | |
| GCG | 88.7 | 18.0 | 14.6 | 11.2 | 44.5 | 11.6 | 9.1 | 2.5 | |
| Input Embed | 92.1 | 16.3 | 13.0 | 15.7 | 80.4 | 16.8 | 12.2 | 9.6 | |
| Average | 80.6 | 19.0 | 14.3 | 11.2 | 32.6 | 9.0 | 6.8 | 3.1 | |
@@
Circuit breaking with Harmfulness Probes (HP). Our proposed circuit breaking method relies on representation control. This section evaluates the potential efficacy of representation reading as an alternative. Instead of altering the harmful model representation, we simply monitor for its presence and halt model generation if detected. We employ the same training dataset (harmful circuit breaker set and retain set) used in the LLM experiments. A linear classifier and an MLP classifier are trained to distinguish between model activations from the two datasets. Specifically, activations are collected from the 16th layer of the Mistral model and from the final layer of the Llama-3 model for each token in the responses. The MLP probe has two layers with hidden size of 64 and 32. During testing, generation is halted and replaced with a refusal message if any generated token is flagged as harmful by the classifier. We find a threshold so that the false positive rate (FPR) on WildChat (Appendix B) is around the same as the model trained with RR [74]. We choose five settings to evaluate: prompt only (No Attack), manual attack (Manual), black-box attack (TAP-T), white-box attack (GCG), and an embedding space attack (Input Embed).
@@
As shown in table Table 2, Harmfulness Probing (HP) significantly reduces the attack success rate compared to the refusal-trained baseline. Both Linear and MLP Harmfulness Probes are outperformed by the representation control approach (RR), however, the gap is smaller for the MLP probe. We emphasize that, although generally probing can be easily thwarted by adversarial attacks, much like input and output filters, its robustness in this context can be largely attributed to the continuous monitoring of model representations associated with harmful processes throughout the entire generation, a key idea in circuit breaking. Furthermore, one can combine HP and RR to implement multiple layers of defense. It is important to note, however, that the Harmful Probes are tested under a weaker adversarial setting, where the attacker lacks knowledge of the probe and does not directly optimize against it. Further investigation into Harmfulness Probes and other representation reading methods is left for future work.
@@ C.4.1 AI Agent Benchmark Example
Appendix D Open LLM Results
| Mistral-7B-Instruct-v2 | Llama-3-8B-Instruct | ||||||
| Benchmarks () | MMLU | 59.1 | 61.3 | 58.9 | 65.6 | 65.0 | 65.6 |
| ARC-c | 62.3 | 60.9 | 62.4 | 62.0 | 61.4 | 63.1 | |
| HellaSwag | 84.8 | 83.0 | 82.6 | 78.6 | 76.8 | 82.6 | |
| TruthfulQA | 66.8 | 45.5 | 67.0 | 51.7 | 51.7 | 60.0 | |
| Winogrande | 76.8 | 78.6 | 77.4 | 75.9 | 76.7 | 78.9 | |
| GSM8k | 42.9 | 38.1 | 44.1 | 78.6 | 78.5 | 81.0 | |
| Average | 65.4 | 61.2 | 65.4 | 68.8 | 68.3 | 71.9 | |
Table 5 shows the scores for each individual benchmark in the Open LLM evaluation.
Appendix E Detailed Results in Multimodal and Agent Settings
@@
Appendix F Multilingual Results
@@
Appendix G Additional Ablation Results
@@
Appendix H Extended analysis of representations